OpenStudy (anonymous):

-4p-10(2p-3)=22-7(1-2p)

OpenStudy (anonymous):

-4p - 20p + 30 = 22 - 7 + 14p -4p + 14p - 20p = 22 - 7 - 30 -10p = 22 - 37 -10p = -15 p = 2/3

OpenStudy (anonymous):

-4p-10(2p-3)=22-7(1-2p) -4p-20p+30=22-7+14p -24p+30=15+14p -34p=-15 p= 15/34

OpenStudy (anonymous):

@pratu043 u got the -14p

OpenStudy (anonymous):

ok we got two differnet answers

OpenStudy (anonymous):

@pratu043 you need to combine similar terms before transpose

OpenStudy (anonymous):

What's wrong with me!! traile is right.

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

-4p-10(2p-3)=22-7(1-2p) 4p-20+30=22-7+14p -24p+30=15+14p -24p-14p=15-30 -39p=-15 p=15/38

OpenStudy (anonymous):

@pastory typo........... it should be p=15/34

OpenStudy (anonymous):

p=15/38